Sarah had a crate of apples. Exactly two-fifths of the apples were red.
Sarah removed 20 red apples from the crate, so that only one-third of the remaining apples were red.
How many apples were left in the crate?

Answer by richwmiller(17219)

2x/5=r
r-20=(x-20)/3
2x/5-20=(x-20)/3
6x/5-60=5/5x-20
x/5=40
x=200 total apples
2/5*200=80 red apples
80-20=60 red apples remain
200-20=180 apples left in the crate
60/180=1/3